OpenStudy (anonymous):

what does f '(1) equal if the line tangent to the graph function f has a point at (1,7) and passes through point (-2,-2)?

OpenStudy (anonymous):

First you must find an equation given two point (1, 7) and (-2, -2). y2-y1/x2-x1 = m m=3 (slope) y=3x+b (gives you general point slope formula (7) = 3(1) + b (plug in a point to find y intercept, b) b=4 gives you the function y=3x+4 f' = 3 therefore f'(1) = 3